OpenStudy (anonymous):

how can i find the least common denominator of 4/9

OpenStudy (anonymous):

4 = 2 * 2 9 = 3 * 3 there are no common factors so the lcd is found by multiply all the factors 2*2*3*3 = 36

OpenStudy (anonymous):

I was not sure if that was the correct answer. I always second guess myself. thanks

OpenStudy (anonymous):

yw